# The metaplectic semigroup and its applications to time-frequency analysis and evolution operators

### v1

@@ 6.1. Covariance and Cohen class
We say that a complex metaplectic Wigner distribution $W_{\mathcal{A}}$ is covariant if for every $z\in{\mathbb{R}^{2d}}$ and every $f,g\in L^{2}(\mathbb{R}^{d})$,
@@
for a suitable kernel $k\in\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$. Because of the convolution nature of the Cohen class, time-frequency representations satisfying (285) are always covariant, see e.g., [17, Proposition 1.3.19 (i)]. On the top of that, it turns out that in the case of metaplectic Wigner distributions the reverse relation also holds:
@@
Let $W_{\mathcal{A}}$ be a complex metaplectic Wigner distribution with projection
@@
up to a constant. The implication $(iii)\Rightarrow(i)$ follows by the fact that every quadratic time-frequency representation in the Cohen class is covariant. It remains to prove that $(i)\Rightarrow(ii)$. Assume that $(i)$ holds. By Theorem 4.6 applied to $f(t)=e^{-\pi|t|^{2}}$ and $(x,\xi)\in{\mathbb{R}^{2d}}$, we have
@@
By writing the blocks (287) of $\mathcal{A}$ explicitly and using (44)–(46), equation (294) is equivalent to $\mathcal{A}$ having blocks (289) with $A_{13},A_{21}\in\operatorname*{Sym}(d,\mathbb{C})$. It remains to impose the condition for $\mathcal{A}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$. By writing $\mathcal{A}=\Re(\mathcal{A})+i\Im(\mathcal{A})=\mathcal{A}_{R}+i\mathcal{A}_{I}$, it is straightforward to see that $\mathcal{A}_{I}^{\top}J\mathcal{A}_{I}=O$, so the positive semi-definiteness condition (104) reads for $\mathcal{A}$ as
(295) $\mathcal{M}_{\mathcal{A}}=\begin{pmatrix}\mathcal{A}_{R}^{\top}J\mathcal{A}_{I}&O_{2d}\\ O_{2d}&\mathcal{A}_{R}^{\top}J\mathcal{A}_{I}\end{pmatrix}\geq 0,$
@@ 6.2. Generalized spectrograms
see [7]. In particular, generalized spectrograms are covariant.
@@ 6.3. Metaplectic Wigner distributions and complex conjugation
Alongside covariance, a fundamental property of the classical Wigner distribution, having implications in the theory of wave fronts, is its behavior with respect to complex conjugation. Indeed, for $f,g\in L^{2}(\mathbb{R}^{d})$
@@
The real case was studied in [30], where it was proven that (327) holds only for the classical Wigner distributions, up to rescalings. This suggests that the classical Wigner distribution is, among metaplectic Wigner distributions with $\widehat{\mathcal{U}}\in\operatorname*{Mp}(2d,\mathbb{R})$, the only quantization tailored to study interaction between time-evolving signals, as it is shown in a recent, unpublished manuscript by two of the authors. In the complex framework, the situation is more involved. Actually, this complication paves the way to many new quantization rules satisfying covariance and (327) within the metaplectic framework.
@@
    $\mathcal{A}\in\operatorname*{Sp}_{+}(2d,\mathbb{R})$ and
@@
By linearity and the density of $\mbox{span}\{f\otimes\overline{g}:f,g\in L^{2}(\mathbb{R}^{d})\}$, it follows that (327) holds if and only if $\widehat{\mathcal{A}}=\widehat{\widetilde{\mathcal{A}}}\mathfrak{T}_{L}$. From the perspective of symplectic projections, this holds if and only if $\mathcal{A}=\widetilde{\mathcal{A}}\mathcal{D}_{L}$. By writing the blocks of $\mathcal{A}$, $\widetilde{\mathcal{A}}$ and $\mathcal{D}_{L}$ explicitly, the equivalence $(i)\Leftrightarrow(ii)$ follows. Let $\widehat{\mathcal{B}}:=\widehat{\mathcal{A}}\widehat{\mathcal{A}}_{1/2}^{-1}$, where $\widehat{\mathcal{A}}_{1/2}$ is the metaplectic operator defining the Wigner distribution, i.e., $W(f,g)=\widehat{\mathcal{A}}_{1/2}(f\otimes\overline{g})$, whose projection $\mathcal{A}_{1/2}$ is given by (226). The equivalence $(ii)\Longleftrightarrow(iii)$ is then a simple restatement of $\widehat{\mathcal{A}}=\widehat{\mathcal{B}}\widehat{\mathcal{A}}_{1/2}$. Since $\mathcal{A}_{1/2}\in\operatorname*{Sp}(2d,\mathbb{R})$, $\mathcal{A}\in\operatorname*{Sp}_{+}(d,\mathbb{C})$ if and only if $\mathcal{B}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$. Moreover, $\mathcal{B}$ is a matrix in the form (257). It follows by Theorem 5.9 that $\mathcal{B}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$ if and only if $\Re(E_{\mathcal{A}})\in\operatorname*{GL}(2d,\mathbb{R})$, $\Re(E_{\mathcal{A}})^{\top}\Im(F_{\mathcal{A}})\geq 0$ and $\Re(E_{\mathcal{A}})(\Im(E_{\mathcal{A}})J)^{\top}\leq 0$. Since $J^{\top}=-J$, the latter condition is equivalent to $\Re(E_{\mathcal{A}})J\Im(E_{\mathcal{A}})^{\top}\geq 0$, and we are done.
@@ 7. Applications to evolution equations with complex quadratic Hamiltonians
The theory of $\operatorname*{Mp}_{+}(d,\mathbb{C})$ allows us to investigate the propagation of concentration for the Wigner distribution associated with solutions of evolution equations of the form
@@ 7.1. The Weyl perspective and modulation spaces
where, for $j=1,\ldots,d$, $a_{j}\in\mathcal{S}^{\prime}(\mathbb{R}^{2})$ is the Weyl symbol of $\widehat{\Xi}_{j}$. If $\widehat{\Xi}_{j}=\mathrm{id}$, then its Weyl symbol is $a_{0,j}(z)=1$. Moreover, the Weyl symbol of $\mathfrak{p}_{i\alpha_{j}}$ is the Gaussian $a_{1,j}(z)=e^{-\pi\alpha_{j}z^{2}}$. Assume $\vartheta_{j}>0$ and $f\in\mathcal{S}(\mathbb{R})$, then by writing the Fourier transform of $f$ in (116), we obtain
@@
(348) $\lVert a\rVert_{M^{\infty,1}_{1\otimes v_{s}}}=\lVert V_{\Phi}a\rVert_{L^{\infty,1}_{1\otimes v_{s}}}=\lVert\zeta\mapsto\lVert V_{\Phi}a\rVert_{\infty}v_{s}(\zeta)\rVert_{1}.$
@@
Let us write $M=(I+V^{-\top}\Sigma V^{-1})^{1/2}$, and $\sigma_{\mathrm{max}}(M)\geq 1$ be the largest eigenvalue of $M$. Then,
@@
Every $\widehat{Z}\in\operatorname*{Mp}_{0}(d,\mathbb{C})$ is a Weyl operator with symbol in $S^{0}_{0,0}({\mathbb{R}^{2d}})$. Consequently, every $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ can be uniquely written as
@@
with $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$, $\mathrm{Op}^{\mathrm{w}}(a)$ Weyl quantization of a real quadratic form $b$, and $c\in S^{0}_{0,0}({\mathbb{R}^{2d}})$, is a one-parameter subgroup of generalized metaplectic operators. The benefits of this perspectives cover different areas of harmonic analyis, with contributions also in applied and computational sciences, due to the characterization of generalized metaplectic operators in terms of their Gabor matrix decay, the infinite dimensional equivalent of sparsity. This approach lies beyond the scope of the present manuscript and will be addressed in future work. The reader may refer to [14, 57] as main references.
@@
where we may assume $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$. The propagator of (361) has been computed in [12], but it is also computable by means of Mehler fomulas. It has polar decomposition
@@ 7.2.1. A prototypical example
where $c(t)\in S_{0,0}^{0}(\mathbb{R}^{4d})$. Therefore, the proof follows as in [18, Theorem 1.6] (cf. [10, Theorem 7.7]).
@@
As a future direction, we stress that Definition 7.9 can be extended to encode the microlocal interactions between two states $f$ and $g$. This was done for metaplectic Wigner distributions $W_{\mathcal{A}}$ with $\widehat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{R})$ in [31]. In particular, by defining the (cross-)Wigner wave front sets $\mathcal{W}\mathcal{F}_{\mathcal{A}}(f,g)$ one can investigate results analogous to those obtained in the aforementioned manusctipt.
@@ 7.2.2. The general case
which yields (394), and once again the proof follows by a microlocalization argument as in [18, Theorem 1.6].

### v2

@@ 6.1. Covariance and Cohen class
We say that a metaplectic Wigner distribution $W_{\mathcal{A}}$ is covariant if for every $z\in{\mathbb{R}^{2d}}$ and every $f,g\in L^{2}(\mathbb{R}^{d})$,
@@
for a suitable kernel $k\in\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$. Because of the convolution nature of the Cohen class, time-frequency representations satisfying (287) are always covariant, see e.g., [16, Proposition 1.3.19 (i)]. On the top of that, it turns out that in the case of metaplectic Wigner distributions the reverse relation also holds:
@@
Let $W_{\mathcal{A}}$ be a metaplectic Wigner distribution with projection
@@
up to a constant. The implication $(iii)\Rightarrow(i)$ follows by the fact that every quadratic time-frequency representation in the Cohen class is covariant. It remains to prove that $(i)\Rightarrow(ii)$. Assume that $(i)$ holds. By Theorem 4.6 applied to $f(y)=e^{-\pi|y|^{2}}$ and $(x,\xi)\in{\mathbb{R}^{2d}}$, we have
@@
By writing the blocks (289) of $\mathcal{A}$ explicitly and using (44)–(46), equation (296) is equivalent to $\mathcal{A}$ having blocks (291) with $A_{13},A_{21}\in\operatorname*{Sym}(d,\mathbb{C})$. It remains to impose the condition for $\mathcal{A}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$. By writing $\mathcal{A}=\Re(\mathcal{A})+i\Im(\mathcal{A})=\mathcal{A}_{R}+i\mathcal{A}_{I}$, it is straightforward to see that $\mathcal{A}_{I}^{\top}J\mathcal{A}_{I}=O_{4d}$, so the positive semi-definiteness condition (103) reads for $\mathcal{A}$ as
(297) $\mathcal{M}_{\mathcal{A}}=\begin{pmatrix}\mathcal{A}_{R}^{\top}J\mathcal{A}_{I}&O_{4d}\\ O_{4d}&\mathcal{A}_{R}^{\top}J\mathcal{A}_{I}\end{pmatrix}\geq 0,$
@@ 6.2. Generalized spectrograms
where $\mathcal{I}F(x)=F(-x)$ is the flip operator, see [7]. In particular, generalized spectrograms are covariant.
@@ 6.3. Metaplectic Wigner distributions and complex conjugation
Alongside covariance, a fundamental property of the classical Wigner distribution, having implications in the theory of Wigner microlocal analysis, is its behavior with respect to complex conjugation. Indeed, for $f,g\in L^{2}(\mathbb{R}^{d})$
@@
The real case was studied in an unpublished contribution  [^3] See G. Giacchi, G., and L. T. Minh. The symplectic Wigner distribution associated to quadratic Fourier transforms: Heisenberg’s uncertainty inequality and applications. 10.36227/techrxiv.176472679.93143839/v1, 2025., where it was proven that (329) holds only for the classical Wigner distributions, up to rescalings. This suggests that the classical Wigner distribution is, among metaplectic Wigner distributions with $\widehat{\mathcal{U}}\in\operatorname*{Mp}(2d,\mathbb{R})$, the only quantization tailored to study interaction between time-evolving signals, as it is shown in a recent, unpublished manuscript by two of the authors  [^4] See G. Giacchi and D. Tramontana. On the microlocal phase-space concentration of Wigner distributions associated with Schrödinger evolutions. arXiv preprint arXiv:2511.19733v2, 2025.. In the complex framework, the situation is more involved. Actually, this complication paves the way to many new quantization rules satisfying covariance and (329) within the metaplectic framework.
@@
    $\mathcal{A}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$ and
@@
By linearity and the density of $\mbox{span}\{f\otimes\overline{g}:f,g\in L^{2}(\mathbb{R}^{d})\}$, it follows that (329) holds if and only if $\widehat{\mathcal{A}}=\widehat{\widetilde{\mathcal{A}}}\mathfrak{T}_{L}$. From the perspective of symplectic projections, this holds if and only if $\mathcal{A}=\widetilde{\mathcal{A}}\mathcal{D}_{L}$. By writing the blocks of $\mathcal{A}$, $\widetilde{\mathcal{A}}$ and $\mathcal{D}_{L}$ explicitly, the equivalence $(i)\Leftrightarrow(ii)$ follows. Let $\widehat{\mathcal{B}}:=\widehat{\mathcal{A}}\widehat{\mathcal{A}}_{1/2}^{-1}$, where $\widehat{\mathcal{A}}_{1/2}$ is the metaplectic operator defining the Wigner distribution, i.e., $W(f,g)=\widehat{\mathcal{A}}_{1/2}(f\otimes\overline{g})$, whose projection $\mathcal{A}_{1/2}$ is given by (228). The equivalence $(ii)\Longleftrightarrow(iii)$ is then a simple restatement of $\widehat{\mathcal{A}}=\widehat{\mathcal{B}}\widehat{\mathcal{A}}_{1/2}$. Since $\mathcal{A}_{1/2}\in\operatorname*{Sp}(2d,\mathbb{R})$, $\mathcal{A}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$ if and only if $\mathcal{B}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$. Moreover, $\mathcal{B}$ is a matrix in the form (259). It follows by Theorem 5.10 that $\mathcal{B}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$ if and only if $\Re(E_{\mathcal{A}})\in\operatorname*{GL}(2d,\mathbb{R})$, $\Re(E_{\mathcal{A}})^{\top}\Im(F_{\mathcal{A}})\geq 0$ and $\Re(E_{\mathcal{A}})(\Im(E_{\mathcal{A}})J)^{\top}\leq 0$. Since $J^{\top}=-J$, the latter condition is equivalent to $\Re(E_{\mathcal{A}})J\Im(E_{\mathcal{A}})^{\top}\geq 0$, and we are done.
@@ 7. Applications to evolution equations with complex quadratic Hamiltonians
Hörmander’s theory has been used to investigate the propagation of concentration for the Wigner distribution associated with solutions of evolution equations of the form
@@ 7.1. The Weyl perspective and modulation spaces
where, for $j=1,\ldots,d$, $a_{j}\in\mathcal{S}^{\prime}(\mathbb{R}^{2})$ is the Weyl symbol of $\widehat{\Xi}_{j}$. If $\widehat{\Xi}_{j}=\mathrm{id}_{L^{2}}$, then its Weyl symbol is $a_{0,j}(z)=1$. Moreover, the Weyl symbol of $\mathfrak{p}_{i\alpha_{j}}$ is the Gaussian $a_{1,j}(z)=e^{-\pi\alpha_{j}z^{2}}$. Assume $\vartheta_{j}>0$ and $f\in\mathcal{S}(\mathbb{R})$, then by writing the Fourier transform of $f$ in (115), we obtain
@@
(350) $\lVert a\rVert_{M^{\infty,1}_{1\otimes v_{s}}}=\lVert V_{\Phi}a\rVert_{L^{\infty,1}_{1\otimes v_{s}}}=\lVert\zeta\mapsto\lVert V_{\Phi}a(\cdot,\zeta)\rVert_{\infty}v_{s}(\zeta)\rVert_{1}.$
@@
Let us write $M=(I+V^{-\top}\Sigma V^{-1})^{1/2}$, and $\sigma_{\mathrm{max}}(M)$ be the largest eigenvalue of $M$. Then,
@@
Every $\widehat{Z}\in\operatorname*{Mp}_{0}(d,\mathbb{C})$ is a Weyl-quantized pseudodifferential operator with symbol in $S^{0}_{0,0}({\mathbb{R}^{2d}})$. Consequently, every $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ can be uniquely written as
@@
with $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$, $\mathrm{Op}^{\mathrm{w}}(b)$ Weyl quantization of a real quadratic form $b$, and $c\in S^{0}_{0,0}({\mathbb{R}^{2d}})$, is a one-parameter subgroup of generalized metaplectic operators. The benefits of this perspectives cover different areas of harmonic analyis, with contributions also in applied and computational sciences, due to the characterization of generalized metaplectic operators in terms of their Gabor matrix decay, the infinite dimensional equivalent of sparsity. This approach lies beyond the scope of the present manuscript and will be addressed in future work. The reader may refer to [13, 55] as main references.
@@
where we may assume $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$. The propagator of (363) has been computed in the unpublished work at Footnote 2, but it is also computable by means of Mehler fomulas. It has polar decomposition
@@ 7.2.1. A prototypical example
where $c(t)\in S_{0,0}^{0}(\mathbb{R}^{4d})$. Therefore, the proof follows as in [17, Theorem 1.6] (cf. [10, Theorem 7.7]).
@@
As a future direction, we stress that Definition 7.9 can be extended to encode the microlocal interactions between two states $f$ and $g$. This was done for metaplectic Wigner distributions $W_{\mathcal{A}}$ with $\widehat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{R})$ in the recent unpublished contribution at Footnote 4. In particular, by defining the (cross-)Wigner wave front sets $\mathcal{W}\mathcal{F}_{\mathcal{A}}(f,g)$ one can investigate results analogous to those obtained in the aforementioned manusctipt.
@@ 7.2.2. The general case
which yields (396), and once again the proof follows by a microlocalization argument as in [17, Theorem 1.6].